OpenStudy (anonymous):

find the y-intercepct for 2x + y=6 Plezz HeLp

OpenStudy (radar):

Substitute 0 for x giving you 6 for y,

OpenStudy (radar):

Intercepts y axis at (0,6)

OpenStudy (anonymous):

thank u so much..

OpenStudy (radar):

NP. Did you understand it?

OpenStudy (anonymous):

would it be 2,0 if not no i didnt

OpenStudy (radar):

Substitute 0 for x in the equation 2x + y = 6 2(0) + y = 6 0 + y = 6 y = 6 Y intercept occurs at (0,6) at x = 0, y = 6

OpenStudy (anonymous):

I got you i didnt kno what happened to the 2 so it would be 2*0=0+y=6 now i see how u got the last number thanks so much for explaining

OpenStudy (radar):

Great, good luck in your studies.